# Imperceptible Protection Against Style Imitation from Diffusion Models

### v1

@@ 3.2 Perceptual Constraint Bank
+ Impasto 0.159 (+0.022) 0.315 (+0.049) 0.912 (+0.016) 4.479 20.74 279.3
AdvDM 0.167 (+0.000) 0.730 (+0.000) 0.846 (+0.000) 3.761 12.45 269.0
+ Impasto 0.136 (+0.031) 0.531 (+0.199) 0.895 (+0.049) 3.897 12.54 271.6
+ Impasto 0.207 (+0.042) 0.709 (+0.073) 0.886 (+0.089) 5.632 10.90 157.8
AdvDM 0.241 (+0.000) 0.776 (+0.000) 0.775 (+0.000) 4.802 10.95 153.5
+ Impasto 0.231 (+0.010) 0.774 (+0.002) 0.797 (+0.022) 4.793 11.87 154.0
Figure 5: Qualitative comparison of baseline methods with and without Impasto. While maintaining comparable style protection efficacy (artifacts in the generated results), Impasto significantly enhances the protected images’ quality.
4 Experiment
Implementation details. We optimize for $N=100$ steps and IWR updates $\mathcal{M}$ for $P=25$ steps after perturbation optimization. Other settings are aligned with that of baseline models; please refer to Suppl. for more details.
Refer to captionRefer to caption
(a) Painting dataset (b) Cartoon dataset
Figure 6: Style protection comparison. Protection performance is evaluated with FID [15] and imperceptibility via DISTS [10]. Adaptation of Impasto to both PhotoGuard [46] and AdvDM [30] ensures superior performance.
Datasets. We utilize two art domain datasets: painting and cartoon. The painting dataset is curated from WikiArt [51] with a selection of 15 artists, 10 works per artist. The cartoon dataset is a collection of 15 cartoons with 10 cartoon face images. Further details are in Suppl.
Evaluation. To evaluate protected image quality, we use DISTS [10], PieAPP [41], and TOPIQ [4]. Protection performance is measured with NIQE [36], BRISQUE [35], and FID [15]. In protection metrics, worse scores indicate more effective protection, as our objective is to prevent style mimicry. It is important to note that some protection assessments, such as NIQE and BRISQUE, are non-reference-based, leading to somewhat inconsistent scores. These tend to fluctuate instead of showing a consistent progression with varying protection strengths. FID, although potentially inconsistent due to a limited number of evaluation samples, aligns more closely with human preferences in this task. Therefore, we further validated Impasto’s effectiveness through human evaluation. For more detailed settings on evaluation metrics, including the user study, please refer to Suppl.
Baseline. Impasto can be integrated into any methods generally formulated in Eq. 4. Based on this, in our benchmark, we incorporate Impasto into PhotoGuard [46] (encoder-based) and AdvDM [30] (UNet-based). In Suppl., applications to Mist [29] and Anti-DreamBooth [52] are also presented.
4.1 Model Comparison
Table 2 presents a quantitative comparison of visual quality under a comparable protection performance of the models with and without Impasto. Across all the scenarios, Impasto substantially enhances the fidelity of the protected images. Fig. 5 also supports the superior efficacy of Impasto; it successfully minimizes artifacts, in contrast to baselines that leave discernible traces on the artwork. It is particularly pronounced in the cartoon dataset (bottom), where both vanilla methods introduce noticeable artifacts in facial areas, potentially disrupting user immersion when reading cartoons. In contrast, Impasto reduces artifacts significantly, rendering them nearly invisible unless examined closely and meticulously. User evaluation further confirms the effectiveness of Impasto (Fig. 7).
Figure 7: User preference study (via A/B test) of PhotoGuard [46] and AdvDM [30] methods with and without Impasto.
Varying protection strengths. In Fig. 6, we manipulate the protection strengths (budgets) to delineate an imperceptibility-protection trade-off curve. On both painting and cartoon datasets, employing Impasto into the baselines (PhotoGuard and AdvDM) considerably improves the trade-off dynamics; Impasto consistently achieves higher imperceptibility with comparable protection performance, or delivers enhanced protection without compromising image quality.
4.2 Model Analysis
Ablation study. In Table 3, we dissect components of Impasto. The proposed PAP markedly improves image fidelity over the base model (Base*), albeit with a slight reduction in protection efficacy. However, compared to the base model with lower protection (Base**), PAP achieves superior preservation of protection performance with enhanced image quality. We observed that when $\mathcal{M}$ is formed from a random mask, not JND, (w/o JND), the protection performance is degraded. One might expect that the JND maintains protection scores while improving image quality. Indeed, the random mask constrains the protection intensity in a similar level to the JND-based map. This is because both maps are normalized between 0 and 1. As a result, they yield similar perturbation magnitudes, leading to a comparable image quality. Nonetheless, the JND-based perceptual map prioritizes less sensitive regions for perturbation while reducing it in highly sensitive areas. Despite the similar protection magnitudes, the JND-based PAP achieves more effective protection. This hints at the possibility that the JND maps are helpful in finding the areas that are important for style protection. For instance, in areas with complex textures, the PAP applies more perturbations than its non-JND counterpart as shown in Fig. 4e. However, these perturbations are typically imperceptible to humans, thereby ensuring perceptually acceptable image quality. Employing a constraint bank significantly improves image quality with comparable protection scores to the baseline (Base**). The omission of the perceptual map $\mathcal{M}$ in the constraints (w/o Mask) leads to a decline in protection performance, akin to the observations in the PAP case; as we confine the influence of constraints in a perceptual manner, it improves protection efficacy.
Table 3: Component analysis. We incrementally attach proposed components to assess their contribution. PAP: perception-aware protection. w/o JND: initialize $\mathbf{M}$ with random masks, not from JNDs. w/o Mask: Constraints without a perceptual map $\mathcal{M}$. Base*: the baseline model (PhotoGuard [46]) with equal protection strength. Base**: the baseline model with a similar level of protection performance.
Method Image Quality Protection Performance
Base* 0.212 0.845 17.66 299.1
+ PAP 0.171 0.890 20.05 286.8
w/o JND 0.170 0.892 20.45 278.1
+ LPIPS 0.163 0.910 21.24 280.4
+ Low-pass 0.163 0.911 21.25 277.5
w/o Mask 0.163 0.911 19.74 272.6
+ CLIP 0.159 0.912 20.74 279.2
Base** 0.181 0.896 20.99 277.6
Table 4: Perception-aware protection. The efficacy of individual JNDs (LA and CM since they show superior results) is presented along with an averaged perceptual map and with IWR. IWR: IWR is conducted before perturbation optimization.
Method Image Quality Protection Performance
Baseline 0.212 0.845 17.66 299.1
LA 0.175 0.879 18.29 284.7
CM 0.169 0.882 17.44 280.8
Average 0.170 0.891 17.42 277.9
IWR 0.171 0.894 19.26 282.5
IWR 0.171 0.890 20.05 286.8
Perceptual map. We also analyze the proposed perception-aware protection in Table 4. LA and CM, which are the best JNDs as demonstrated in Table 1, enhance the fidelity of the protected images but at the cost of significantly compromising protection performance. Creating a perceptual map with multiple JNDs through averaging leads to better image quality compared to scenarios with a single JND. However, this approach also substantially reduces the effectiveness of protection. We speculate that such a straightforward averaging method may not adequately capture the unique structural elements of the image, resulting in overly smooth perturbations that could weaken the protection performance. On the other hand, IWR enhances all protection scores while preserving satisfactory image quality, as it can adapt to the specific textures and structures of a given artwork. It’s noteworthy that applying IWR prior to perturbation optimization (IWR), where it does not consider the perturbations, slightly diminishes protection performance, accentuating the importance of modeling the interplay between artwork and applied perturbation to finalization mask $\mathcal{M}$.
Refer to captionRefer to caption
(a) Against personalization methods    (b) Against diffusion models
Figure 8: Generalization. Impasto does not impede baselines’ generalization abilities on (a) diverse personalization methods; DreamBooth [44], LoRA [17], and DreamStyler [1] or (b) model black-box scenario; optimize on SD v1.5 and test on SD v2.1.
Figure 9: Evaluation on robustness. Methods with Impasto exhibit comparable protection performance to baselines.
Countermeasures. To analyze the robustness of Impasto, we conduct countermeasure experiments involving JPEG compression ($q=40$), Gaussian blur ($3\times 3$ kernel, $\sigma=0.02$), and Gaussian noise ($\sigma=0.02$). Results indicate a performance degradation of all protection methods when these countermeasures are applied, as they tend to remove the protective perturbations (Fig. 9). Nonetheless, Impasto demonstrates comparable robustness against such countermeasures. For the blur effect, there is a slight performance degradation with Impasto, as subtle perturbations are particularly weakened to this; we also observed that baseline with low-budget protection is also vulnerable to blur operation. However, the blurring effect also renders the artwork less plausible, making it an impractical choice for malicious users. The most commonly used method is likely JPEG compression, as it preserves the artwork’s fidelity while being readily applicable.
Black-box scenario. We extend to evaluating the impact of Impasto on the performance of protection when trained with other personalization methods. We adopt LoRA [17], a standard personalization technique in the art creation community, and DreamStyler [1], known for its effectiveness in style adaptation using a textual inversion approach. As illustrated in Fig. 8a, Impasto maintains robust protection effectiveness even with these personalization methods.
In style protection, generalization robustness against unknown models is also a crucial aspect. To examine this, we compare by optimizing on SD v1.5 and testing on SD v2.1, following the setups of Van et al. [52]. As illustrated in Fig. 8b, Impasto successfully preserves the robustness of PhotoGuard and AdvDM. Overall, in both facets of the black-box scenario; unknown personalization methods and diffusion models, Impasto not only maintains the generalization ability but also effectively reduces the artifact on the protected images.
@@ 4.2 Model Analysis
4.3 Impasto in Other Domain
Although Impasto is initially proposed to prevent style imitation, its applicability can be extended beyond other domains or applications. To validate this, we conduct protection on two facial datasets, CelebA-HQ [20] and VGGFace2 [3]. As demonstrated in Table 5 and Fig. 10, adopting Impasto in these natural domains can also enhance the quality of protected images without compromising the protection performance of baseline models. This broad applicability highlights Impasto’s versatility and potential as a universal tool for protecting users’ copyright and preventing serious threats of deepfakes.
5 Conclusion
We have introduced Impasto to prevent style imitation in perceptual orientation. With a proposed perceptual map, Impasto markedly improves the quality of the protected images. A perceptual constraint bank further boosts performance, establishing our method as a versatile and superior protector of artwork.
Limitations. Current protection methods mostly adopt adversarial perturbations, which can hamper usability due to the extensive time required for the optimization. Even accessible software [49] takes 30-60 mins to protect a $512^{2}$ image on an M1 Max CPU. Addressing this time constraint is a challenge that future research should aim to overcome.
Appendix 0.A Method Details
In this section, we describe the details of Impasto— the computation and discussion of JND estimations (Sec. 0.A.1), perceptual constraints (Sec. 0.A.2), and the implementation specifics (Sec. 0.A.3).
0.A.1 JND Estimations
In our study, we investigated five JND models, and here we we provide a comprehensive elucidation of the fundamental rationale of these estimations and delve into the specifics of their computations:
Luminance adaptation (LA): The visibility thresholds within the human visual system vary to luminance levels [19]. For instance, our eye is more sensitive under moderate lighting conditions, while discrimination is challenging in a very dark light. Similarly, we observed that perturbations are less noticeable in regions with extremely low or high luminance, prompting an increase in protection strength within these areas. To implement this concept, luminance adaptation is computed following Chou and Li [5] as in the below equation.
Contrast masking (CM): It is obvious that stimuli become less perceptible against patterned, non-uniform backgrounds. In our context, perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this effect, we first compute the luminance contrast [26] by convolving an input image with four directional filters. This process highlights the complexity of a region by contrasting it with its immediate surroundings. Subsequently, contrast masking is established by mapping the luminance contrast onto a logarithmic curve [57].
where luminance contrast $LC(\mathbf{x})$ is obtained via four directional filters as $LC(\mathbf{x})=\frac{1}{16}\times\text{max}_{k=1,\dots,4}|\mathbf{x}*\nabla_{k}|$ following Wu et al. [58].
Contrast sensitivity function (CSF): The human visual system exhibits a band-pass response to spatial frequency and CSF represents a function of how our eye is sensitive to the contrast of signals at various spatial frequencies. High spatial frequencies correspond to rapid changes in image details, and contrast sensitivity is optimal at moderate spatial frequencies. Beyond a certain frequency threshold, known as the resolution limit, our eyes are unable to detect changes. Upon this concept, several studies have introduced and modified the CSF [25, 34, 8, 33]; the CSF model $H(f,\theta)$ is given below.
$H=\begin{cases}2.6\times(0.0192+0.114f_{\theta})e^{-(0.114f_{\theta})^{1.1}},&\text{if }f\geq 7.8909\\ 0.981,&\text{otherwise.}\end{cases}$ (14)
Here, $f$ is the radial spatial frequency, measured in cycles per degree of visual angle (c/deg). The variable $\theta$, ranging from $[-\pi,\pi]$, denotes the orientation. Additionally, $f_{\theta}=f/[0.15\times\text{cos}(4\theta)+0.85]$ accounts for the oblique effect [25].
The CSF model is now applied to an image in the frequency domain as $\mathbf{x}_{csf}=\mathcal{F}^{-1}[H(u,v)\times\mathcal{F}(\check{\mathbf{x}})]$, where $\mathcal{F}[\cdot]$ and $\mathcal{F}^{-1}[\cdot]$ represent the DFT and its inverse, respectively. Here, $H(u,v)$ is the DFT-version of $H(f,\theta)$, with $u,v$ being the DFT indices. The transformation from $H(f,\theta)$ to $H(u,v)$ is elaborated in Larson et al. [25]. Prior to applying the CSF model, an input image $\mathbf{x}$ is converted into a perception-adjusted luminance form to reflect the non-linear relationship between digital pixel values and physical luminance [25]. This is achieved by calibrating $\mathbf{x}$ to the settings of an sRGB display and converting it into perceived luminances, which indicates the relative lightness: $\check{\mathbf{x}}=\sqrt[3]{(0.02874\mathbf{x})^{2.2}}$.
Standard deviation: It has served as a crucial metric for quantifying the structural information of an image, both in pixel space [55] and feature space [10]. Given that perturbations tend to be less noticeable in areas exhibiting high levels of change, (as discussed in contrast masking), we focus on these sudden changes, interpreting them as structural components. For this purpose, we compute the block-wise standard deviation in pixel space using a $9\times 9$ local block.
Entropy: Similar to standard deviation, block-wise entropy is calculated using a $9\times 9$ local window. This approach is based on the concept that the complexity of a local region influences the detectability of subtle perturbations. The higher the local entropy, the more intricate the region, potentially rendering minor perturbations less perceptible.
Post-processing: All the JND estimations are min-max normalized and then inverted by subtracting from one. When constructing a perceptual map $\mathcal{M}$ with JND estimations, we encountered issues with some JNDs displaying extremely skewed distributions. Additionally, the distributions of JNDs often do not align with each other, posing challenges in their combination. While standardization aligns the distributions, we found that the continuous JND values provide weak signals as masks (e.g. on average, the protection strength drops to 50% compared to the original). To give a more distinct signal depending on the JND values, we discretize the scores by quantizing the JND values. This involves calculating JND quantiles, where the first quantile is set to 1.0. For each subsequent quantile, we multiply by a factor of $\beta$ to decrease the value in a discrete manner. We use $\beta=0.85$, resulting in quantized JND values across four quantiles as [1.0, 0.85, 0.7225, 0.6141]. This allows for a more nuanced adjustment of the protection intensity while accommodating the varied distributions of JND estimations.
0.A.2 Perceptual Constraints
Masked LPIPS. We use AlexNet [23] as the backbone network for the LPIPS loss, adhering to the parameter settings in the official LPIPS documentation.
Masked low-pass. In this constraint, we utilize $\text{LP}(\cdot)$, a reconstruction function focusing on the low-frequency component. Inspired by Luo et al. [31], we implement a DWT-based reconstruction module. Given an input image $\mathbf{x}$, DWT decomposes it into one low-frequency component and three high-frequency components, as expressed by the following equations.
$\mathbf{x}_{ll}=\mathbf{LxL}^{T},\;\;\mathbf{x}_{lh}=\mathbf{HxL}^{T},\;\;\mathbf{x}_{hl}=\mathbf{LxH}^{T},\;\;\mathbf{x}_{hh}=\mathbf{HxH}^{T}.$ (15)
0.A.3 Implementation Details
When we implement Impasto into the existing protection methods, we adhere to their respective settings for optimizing perturbations. Regarding the Impasto’s own components, we use following hyperparameters depicted in Eq. 11 as: $\lambda_{L}=5.0,\lambda_{LP}=10.0,\lambda_{C}=0.1$. For our PAP (Eq. 6), we utilize the $L_{\infty}$ norm. On the other hand, for the IWR loss (Eq. 7), we employ the $L_{2}$ norm. When we calculate IWR loss, since the distance between $\mathcal{L_{SP}}$ respect to $\mathcal{M}^{\prime}$ and $\mathcal{M}(\omega)$ (first term) is significantly smaller than the $L_{2}$ constraint (second term), we amplify the former term by a factor of $5\times 10^{7}$. In addition, during our experiments, it was observed that the magnitudes of $\mathcal{L_{E}}$ and $\mathcal{L_{SD}}$ differ considerably, with the SD loss exhibiting much smaller values. Therefore, when integrating Impasto into UNet-based protection methods (e.g. AdvDM [30], Anti-DreamBooth [52], Mist [29]), we further scale down all the $\lambda_{L},\lambda_{LP},\lambda_{C}$ by a factor of 0.05. This adjustment is made to balance the influence of our proposed components across both protection methods, ensuring that the impact of Impasto is consistent and effective in enhancing image protection.
Appendix 0.B Experimental Settings
Datasets. To create the “painting” dataset, we gather artworks from the WikiArt dataset [51]. This dataset includes selections from 15 artists, 10 artworks per each. The “cartoon” dataset is sourced from the NAVER WEBTOON platform; images are cropped to contain only character depictions. This comprises 15 cartoons with 10 cartoon character images each.
Evaluation. To assess the quality of the protected images, we utilize three deep learning-based full-reference quality assessments: DISTS [10], PieAPP [41], and TOPIQ [4]. These measures compare the protected images against the original artworks. DISTS is designed to be sensitive to structural changes while exhibiting explicit tolerance to texture resampling. It achieves this by evaluating both the mean and correlation of feature maps from the compared images. PieAPP measures the perceptual error of a distorted image with respect to a reference and its training dataset. TOPIQ adopts a top-down approach, leveraging high-level semantics to direct the quality assessment network’s focus towards semantically significant local distortion regions.
@@ Appendix 0.B Experimental Settings
For evaluating protection performance, we employ NIQE [36], BRISQUE [35], and FID [15]. NIQE and BRISQUE are blind image quality assessment methods that only use generated images for evaluation. BRISQUE operates on the premise that distortions in natural images disrupt pixel distributions as well. It processes an input image and then extracts features, which are finally mapped to a mean opinion score using an SVM regressor [47]. NIQE, in contrast, does not rely on opinion scores but instead assesses the quality of an image by comparing its statistical properties with those of a clean image dataset. Despite their widespread use, these methods have limitations in our context: 1) They are designed primarily for natural image domains, making their measurements potentially unreliable in artistic domains, particularly for cartoons. 2) Traditional non-reference assessments focus on common distortions such as JPEG compression, blurring, and Gaussian noise, which may not effectively capture the unique artifacts in generated images in our task. FID, conversely, measures the distance between two distributions—the original artwork and the generated image—providing more reliable assessments. For example, as protection strength increases, FID is the only metric where scores worsen.
@@
User study. We conducted a user study in the form of an A/B test with a reference image as a benchmark. A total of 44 participants were involved, tasked with determining the better-quality protected image (given the reference original artwork) and identifying the lower-quality generated image (given a reference DreamBooth generated image). Each participant was asked to vote on 8 questions for PhotoGuard [46] and another 9 for AdvDM [30].
Appendix 0.C Additional Analyses and Results

### v2

@@ Appendix C Experimental Settings
Painting PhotoGuard 0.181 0.364 0.896 4.306 (+0.000) 20.99 (+0.00) 277.6 (+0.0)
+ Impasto 0.186 0.346 0.889 4.451 (+0.145) 22.02 (+1.03) 286.9 (+9.3)
AdvDM 0.167 0.730 0.846 3.761 (+0.000) 12.45 (+0.00) 269.0 (+0.0)
+ Impasto 0.164 0.503 0.842 3.947 (+0.231) 11.60 ( - 0.85) 276.6 (+7.6)
Cartoon PhotoGuard 0.249 0.782 0.797 5.037 (+0.000) 10.19 (+0.00) 155.9 (+0.0)
+ Impasto 0.241 0.894 0.779 5.480 (+0.443) 12.88 (+2.69) 162.5 (+6.6)
AdvDM 0.241 0.776 0.775 4.802 (+0.000) 10.95 (+0.00) 153.5 (+0.0)
+ Impasto 0.243 0.819 0.744 4.780 ( - 0.022) 11.90 (+0.95) 155.9 (+2.4)
Original artwork No protection Partial protection Full protection
Figure 11: Comparison of restriction types. Extending Figure 2, we compare the fine-tuning results of the original artwork with no protection, partial protection, and full protection. The red inset in the original artwork highlights the area where perturbations were applied in the partial protection case, with perturbations confined to the inset.
Evaluation. To assess the quality of the protected images, we utilize three deep learning-based full-reference quality assessments: DISTS (Ding et al. 2020), PieAPP (Prashnani et al. 2018), and TOPIQ (Chen et al. 2023). These measures compare the protected images against the original artworks. DISTS is designed to be sensitive to structural changes while exhibiting explicit tolerance to texture resampling. It achieves this by evaluating both the mean and correlation of feature maps from the compared images. PieAPP measures the perceptual error of a distorted image with respect to a reference and its training dataset. TOPIQ adopts a top-down approach, leveraging high-level semantics to direct the quality assessment network’s focus towards semantically significant local distortion regions.
To evaluate protection performance, we employ NIQE (Mittal, Soundararajan, and Bovik 2012), BRISQUE (Mittal, Moorthy, and Bovik 2011), and FID (Heusel et al. 2017). NIQE and BRISQUE are blind image quality assessment methods that only use generated images for evaluation. BRISQUE operates on the premise that distortions in natural images disrupt pixel distributions. It processes an input image and then extracts features, which are finally mapped to a mean opinion score with SVM regressor (Schölkopf et al. 2000). NIQE, in contrast, does not rely on opinion scores, instead assesses the quality of an image by comparing its statistical properties with those of a clean image dataset.
@@
It is important to note that some protection assessments, such as NIQE and BRISQUE, despite their widespread use, have limitations in our context: 1) They are designed primarily for natural image domains, making their measurements potentially unreliable in artistic domains, particularly for cartoons. 2) Traditional non-reference assessments focus on common distortions such as JPEG compression, blurring, and Gaussian noise, which may not effectively capture the unique artifacts in generated images in our task. Therefore, these metrics tend to fluctuate instead of showing a consistent progression with varying protection strengths. FID, although potentially inconsistent due to a limited number of evaluation samples, aligns more closely with human preferences in this task. We suspect that it because FID measures the distance between two distributions—the original artwork and the generated image- providing more reliable assessments. Therefore, we further validated Impasto’s effectiveness through human evaluation.
@@
User study. We conducted a user study in the form of an A/B test with a reference image as a benchmark. A total of 44 participants were involved, tasked with determining the better-quality protected image (given the reference original artwork) and identifying the lower-quality generated image (given a reference DreamBooth (Ruiz et al. 2023) generated image). Each participant was asked to vote on 8 questions for PhotoGuard (Salman et al. 2023) and another 9 for AdvDM (Liang et al. 2023).
Appendix D Discussions and Analyses
Full vs. partial protection. Extending Figure 2, in Figure 11, we compare the original artwork, and the personalization results of unprotected, partially protected, and fully protected artworks. Partial protection simulates hard restrictions, whereas full protection mimics soft restrictions. To create partially protected images, we apply protection to the center (Figure 11, top) or the facial region (Figure 11, bottom), as indicated by the red boxes. Both the unprotected and partially protected images cannot prevent against style imitation from the personalization method. In contrast, full protection (soft restriction) successfully prevents this.
We also conduct a quantitative comparison between partial and full protection with FID on the painting dataset. In this experiment, the partial protection achieves 260.2, while full protection scores 299.1, demonstrating the higher efficacy of full protection. This empirical evidence led us to propose soft restriction-based JND perceptual protection.
JNDs. In this section, we further analyze the protection performance of JND-based perception map depicted in Figure 3 (right). When compared to the baseline, perceptual map improves image quality (DISTS) across the board, but it also leads to a compromise in protection performance (FID). Among the JNDs, LA demonstrates the best protection performance. We conjecture that in many artworks, the majority of areas fall high or low-luminance, thereby maintaining perturbations strength high across extensive regions. However, as shown in upper image, even simple textures (e.g. sky) can have strong perturbations, placing LA at the lower image quality. On the other hand, CSF, Std, and Entropy generally apply high perturbations only to specific areas, such as edges, resulting in most regions being not fully protected and consequently, causing a huge degradation in protection performance. CM’s protection intensity is also determined by spatial changes but this covers more detailed local regions (see fields in upper image) and also being based on contrast, leading to both high quality and effective protection.
Ablation study. In Table 2, one might expect that the JND (PAP with JND) maintains protection scores while improving image quality. Indeed, the random mask (PAP without JND) constrains the protection intensity in a similar level to the JND-based map. This is because both maps are normalized between 0 and 1. As a result, they yield similar perturbation magnitudes, leading to a comparable image quality. Nonetheless, the JND-based perceptual map prioritizes less sensitive regions for perturbation while reducing it in highly sensitive areas. Despite the similar protection magnitudes, the JND-based PAP achieves more effective protection. This hints at the possibility that the JND maps are helpful in finding the areas that are important for style protection. For instance, in areas with complex textures, the PAP applies more perturbations than its non-JND counterpart. However, these perturbations are typically imperceptible to humans, thereby ensuring perceptually acceptable image quality.